How much more would you need to deposit at the end of every month to accumulate $288,500 over 20 years if the interest rate earned is 5.0% compounded annually instead of 5.0% compounded monthly?
Answer #1

If interest rate compounded annually:
Rate = 5% / 12
FV = 288,500
Nper = 20 * 12 = 240
PV = 0

Monthly deposit can be calculated by using the following excel formula:
=PMT(5%/12,240,0,-288500)
= $701.89

If interest rate compounded monthly:
Effective rate = (1 + 5% / 12)^12 - 1 = 5.1162%

Monthly rate = 5.1162% / 12
FV = 288,500
Nper = 20 * 12 = 240
PV = 0

Monthly deposit can be calculated by using the following excel formula:
=PMT(5.1162%/12,240,0,-288500)
= $692.52

How much more you need to deposit = $701.89 - $692.52 = $9.37